OpenStudy (anonymous):

HELP ME PLEASE. one acetic acid solution is 70% water, and another 30% water. how many liters of each solution should be mixed to produce 20 liters of a solution that is 40% water? solve this by using a system of equation.

OpenStudy (welshfella):

I misread the question...

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{u have two equations}\\~\\ \dfrac{70-40}{40-30}&=\dfrac{y}{x}........(1) \\~\\ x+y&=20............(2) \end{align}}\)

OpenStudy (mathmath333):

http://www.purplemath.com/modules/mixture.htm

OpenStudy (radar):

20 liters of 40% water is 8 gal of water. You will need 8 gal of water to obtain this mixture. This 8 gal of water is going to come from the 2 mixtures of which one is 70% water and the other is 30% water.

OpenStudy (radar):

Let x equal the amount of 70% water solution. 20 - x will then equal the amount of the other solution (30% water) then the actual water from each of these solutions would be (as given by the problem) .7x + .3(20 - x) = 8 .7x + 6 - .3x = 8

OpenStudy (radar):

Solve for x Solve for 20 - x

OpenStudy (radar):

@katchen_97 Did you follow with understanding?